OpenStudy (anonymous):

a rectangle's width is 5 feet less than its length. if the area of the rectangle is 546 square feet, what is its length, in feet?

OpenStudy (anonymous):



OpenStudy (anonymous):

546 = x(x-5) x^2 - 5x = 546 x^2 - 5x - 546 = 0 Solved: http://prntscr.com/2x4wsv

OpenStudy (anonymous):

so the final answer is 0?

OpenStudy (anonymous):

No, you set the expression equal to zero and solve for x. You have to factor. Click the link http://prntscr.com/2x4wsv for the answer.

OpenStudy (anonymous):

thank you!